OpenStudy (anonymous):

simplify 5 sqrt 3+ 8 sqrt 3 - 6 sqrt 2 + 4 sqrt 2

OpenStudy (anonymous):

\[5\sqrt{3}+8\sqrt{3}-6\sqrt{2}+4\sqrt{2}\] =\[13\sqrt{3}-2\sqrt{2}\]